OpenStudy (anonymous):

Simplify the expression. (4)^-2

OpenStudy (anonymous):

flip it upside down

OpenStudy (anonymous):

It's not a fraction

OpenStudy (anonymous):

yet there is no such thing as a negative exponent

OpenStudy (anonymous):

its -4*-4

OpenStudy (anonymous):

so it would equal..... positive 16

OpenStudy (anonymous):

i16

OpenStudy (anonymous):

Thanks...appreciate all your help

jhonyy9 (jhonyy9):

you need using formula (a)^(-x) = 1/a^(x)

OpenStudy (anonymous):



OpenStudy (anonymous):

When you see negative in exponent flip every inside the parenthesis So, Let's rewrite (4)^(-2) into (4/1)^(-2) Flip in order to turn the exponent into positive (1/4)^(2) = (1/16)